OpenStudy (anonymous):

6x^2+19x+15 factor the expression

OpenStudy (unklerhaukus):

this one is a bit trickier, you might want to use the quadratic formula The Quadratic formula\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[(x-x_1)(x-x_2)=0\]

OpenStudy (unklerhaukus):

so first, identify the coefficients a,b,c